OpenStudy (anonymous):

Two families paid a total of $44.00 to see a movie. Each adult ticket cost $8.00. If there were 3 adults and 4 children, how much did each child's ticket cost? $5.00 $5.50 $6.00 $6.50

OpenStudy (anonymous):

$5.00

OpenStudy (anonymous):

cost of 3 adult ticket = 8x3 = $24 So cost of 4 child tickets = 44-24 = 20 So cost of one child ticket = 20/4 = $5

OpenStudy (anonymous):

Thanks! :) Awesome answer.

OpenStudy (anonymous):

u r welcome .....:))